OpenStudy (anonymous):

Write an equation of the line through (–4, 4) and (8, 8) in point-slope form.

hero (hero):

Hint: Use the slope formula to find the slope of the line.

hero (hero):

Sorry my computer has been lagging badly

OpenStudy (anonymous):

nah its cool

OpenStudy (anonymous):

the slope is 1/3 where do i go from here?

OpenStudy (anonymous):

@jim_thompson5910 do you know?

hero (hero):

Now you use this formula \( y - y_1 = m(x - x_1)\) substitute m = 1/3, y_1 = 4, x_1 = -4

hero (hero):

Then isolate y afterwards and simplify

OpenStudy (anonymous):

is y-4=1/3(x-4) the right answer

hero (hero):

is y isolated?

hero (hero):

Don't you have to add 4 to both sides to isolate y? Then afterwards, it still must be in the form y = mx + b because that's the point slope form.

OpenStudy (anonymous):

wait so what answer did u get cuz none of the multiple choice answers look like that

OpenStudy (anonymous):

please help @Hero